How would you compute the impeadence of a inductor and capictor if the real parts of the complex number were not given.

so instead of it being 4+j6 how would to solve just j6. would you have to have a real number?
Answer #1

Impedance is the combination of resistance and reactance.

if the resistance value is zero then total impedance becomes reactance.

inductive reactance= w*L

capacitive reactance=1/(w*C)

where w= angular frequency =2*3.14*frequency.

The frequency used in US is 50HZ.

To calculate impedance of inductor or capacitor there is no need of resistance part.